OpenStudy (anonymous):

Solve the recurrence relation:

OpenStudy (anonymous):

\[ a_{n} = a_{n-1} + n, a_{0} = 1\]

OpenStudy (loser66):

solve what? just plug n =1, 2, 3.... in.

OpenStudy (loser66):

oh, solve for the formula, right? wow... it was 3 years ago when I took this course. hehehe...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

glad you still got it

OpenStudy (anonymous):

i know they come out to be 2, 4, 7 but I don't see no pattern there with exponents or anything..... ugh

OpenStudy (anonymous):

no your reply!

OpenStudy (loser66):

\[a_1 = a_0+1\\a_2= a_1+2= a_0+1+2=a_0+3\\a_3= a_2+3= a_0+3+3=a_0+6\\.......\\a_n=???\]

OpenStudy (loser66):

All you have to do is interpret the number at the end of each row. You see, at row 2, that is \(a_\color{red}{2}\), then the last number is 3 row 3, \(a_\color{blue}{3}\) then the last number is 6 do some more, you can construct the link between them. Note: the first term is \(a_0\) always, no need to look at it.

OpenStudy (anonymous):

then \(a_{4} = a_{0} + 10 \) and i don't see a link

OpenStudy (loser66):